find the slope from these coordinates
(5,0) (-6,-4)

Answer by checkley77(12844)

(5,0) (-6,-4)
Slope=(y2-y1)/(x2-x1)
Slope=(-4-0)/(-6-5)=-4/-11=4/11 ans.